# An electron-hadron collider at the high-luminosity LHC

### v1

@@ Abstract
We discuss a concept of a lower-energy version of the Large Hadron-electron Collider (LHeC), delivering electron-hadron collisions concurrently to the hadron-hadron collisions at the high-luminosity LHC at CERN. Assuming the use of a 20 GeV electron Energy Recovery Linac (ERL), we describe the optimised beam dynamics, accelerator technologies, and detector constraints required for such a ”phase-one” LHeC. Finally, we also discuss the ERL configurations, the possibility of delivering electron-hadron collisions during the planned LHC Run5 and briefly outline the scientific potential of this proposal.
@@ 1 Introduction
The Large Hadron-electron Collider (LHeC) [1] will open new frontiers in high-energy physics as it will become a unique laboratory for studies of deep-inelastic scattering (DIS) and Higgs boson, electroweak, top quark and BSM phenomena in a vast range of virtualities $Q^{2}$, from very close to zero up to $10^{6}$ GeV2, and for a much extended range of $x$ – the fractional longitudinal momentum of partons. An intense electron beam from a novel energy recovery linac (ERL) will collide with a proton or ion beam from the high-luminosity LHC (HL-LHC), in the re-designed P2 interaction region capable of hosting either electron-hadron or hadron-hadron collisions, concurrently with hadron-hadron collisions in other HL-LHC experiments [2].
In the Conceptual Design Report (CDR) of LHeC, it is assumed that a three-pass ERL delivers a 50 GeV high-current electron beam [1]. In this proposal, we assume the use of a 20 GeV single-pass ERL (see Fig. 1) with yet higher beam current of 60 mA, and with a circumference of a third of the LHC’s. This choice of design is motivated by the availability of the technology to operate a single-pass ERL (although presently at lower beam currents) and by the overall simplification, in particular, in the design of the adequate machine-detector interface thanks to much lower levels of synchrotron radiation. One should note that the cost of building such a ”phase-one” LHeC as well as its running costs are significantly lower than for the final LHeC configuration.
Figure 1: Schematics of the single-pass LHeC layout with the optional return arcs $3-6$ (dashed lines).
Such a staged realisation of the LHeC project provides a huge benefit not only in reducing financial strains but first of all in the availability of concurrent electron-hadron collisions already during Run5 of the HL-LHC. Consequently, apart from the possibility of performing a wide range of unique and important studies almost 10 years earlier, the results of such a new electron-hadron experiment could also be used to timely improve measurements in other (hadron-hadron) experiments at the HL-LHC.
@@ 2 The LHeC interaction region
The fundamental principle of an ERL provides an ideal condition for a staged particle accelerator. Still, a number of items have to be followed up, as beam parameters scale with energy, and so while keeping the general layout untouched, a re-optimisation of the beam dynamics is needed for highest performance of the electron-proton collisions.
Schematically, the concept is shown in Fig. 1 – the geometry of the ERL will remain the same: two superconducting linacs are connected by return arcs on both sides. Details about the general layout can be found in [3] and [4]. However, for the phase-one case with reduced electron energy modules needed to reach higher energies are left out: arcs 3 to 6 will not be installed, nor will the so-called bypass needed in the final 50 GeV design to guide the beam outside the particle detector, which will only be installed in a later staging step. Beam spreaders and re-combiner modules will not be present in the first-stage lattice.
On the other hand, a number of key elements will define the geometry of the ERL from phase-one and remain untouched for all energy steps. The linacs, the interaction region (IR) and the beam separation scheme will be designed to be valid for any operation energy. As described in detail in [3] the proton and electron beams are separated after the interaction point through their different beam rigidities. A combination of a weak dipole field embedded inside the detector and the combined function effect of the off-centred mini-beta quadrupoles of the electron lattice acts as an effective separation scheme. The actual value of the dipole strength and the offset of the quadrupoles were optimised for the lowest synchrotron radiation and thus equal effective bending radii in the three magnets [5], and for the phase-one scenario described here, the electron bending radius of this separation scheme is of about 1500 m.
@@
  • $p$-optics, orbit / beta-function
  • beam-beam effect on the proton beam / ditto for the electrons
  • electron emittance
  • luminosity reach
  • beam separation scheme and synchrotron radiation
  • polarisation
  • cost savings
The interaction region of the $ep$ collisions is shown schematically in Fig. 2.
Figure 2: The interaction region of electron and proton beams: in black the mini-beta focusing scheme of the electrons is shown, including the detector dipole field that is part of the beam separation. Both are embedded within the free space of the HL-LHC proton inner triplet lattice, marked in blue and red.
$p$-optics, orbit / beta-function. While the electrons will be separated early enough and do not see any influence from the proton mini-beta quadrupoles, the proton beam on the other side will be distorted by the fields of the electron mini-beta quadrupoles. However, the influence on beam optics and orbit is weak and scales with the ratio of the energy of both counter-rotating beams. For the design case of the 50 GeV electron and 7 TeV proton beams, the effect has been calculated and the remaining distortions compensated; see Fig. 3. The so-called beta-beat, $\Delta\beta/\beta$ that reaches not more than 1.8% is locally corrected by a rematch of the neighbouring quadrupoles [3]. Reducing the energy of the electron beam to 20 GeV evidently also reduces this effect, and values of the order of 0.7% are expected at most. Compared to the overall budget of $\Delta\beta/\beta=20\%$ for the LHC optics tolerance this is a negligible contribution; nevertheless, a correction is available and has been applied.
Figure 3: Local distortion of the proton optics due to the influence of the electron quadrupoles before (red) and after a local compensation of the effect (blue).
@@
where we refer here to the effect on the proton beam; accordingly $N_{e}$ is the electron bunch population, $r_{0}$ the classical proton radius, $\gamma$ is the electron Lorentz factor and $\beta^{\ast}$ and $\sigma$ are the amplitude function and beam size at the Interaction Point (IP). We would like to point out that in this approximation the tune-shift does not depend on the $\beta$-function. However, it depends on the energy of the beam considered. So in case of the protons, the resulting effect caused by the electron beam leads to a tune-shift of $\Delta Q_{x,y}\approx 3.8\cdot 10^{-4}$, nearly two orders of magnitude smaller than the effect of the proton-proton interactions for the HL-LHC operation where we obtain $\Delta Q_{x,y}\approx-8.6\cdot 10^{-3}$ per IP and thus the additional distortion due to $ep$ collisions only has a marginal effect.
@@
Beam-beam effect on the electron beam. However, the situation changes considerably for the electrons. Due to the lower beam energy the electrons are more sensitive, and especially the effect of the non-linear terms has to be minimised. In Fig. 4 the beam size of the electron beam is shown as a function of the position $s$ along the beam orbit with $s=0$ referring to the IP. In black, the theoretical beam size is plotted as it is obtained without beam-beam forces. The strong additional focusing due to the space-charge forces of the proton bunch is shown in blue. Here we would like to point out that the dominant parameter, the proton bunch population, refers to the ultimate goal of $N_{p}=2.2\cdot 10^{11}$ as foreseen for the HL-LHC luminosity upgrade. Under the influence of this additional focusing – called “pinch effect” – the electron beam reaches a smaller effective beam size which is however to a large extend compensated by the so-called hourglass effect of the beam focusing in the vicinity of the IR. The optics distortion however has to be corrected in order to obtain well-matched beam parameters before entering the return arc and the decelerating process of the ERL. This is also highlighted in Fig. 4, where black and blue curves are indeed congruent before and after the interaction region.
@@
Figure 4: Effect of the space charge force on the 20 GeV electron beam: The beam size is reduced (leading to an additional enhancement of the luminosity, pinch effect) and re-matched to the ideal conditions before entering the return arc for energy recovery.
@@
While this fact represents the prerequisite for a successful energy recovery, the non-linear terms of the beam-beam interaction are better visible in the phase-space representation. The situation for the two electron energies considered here, 50 GeV and 20 GeV, is compared in Fig. 5. As expected, the beam-beam effects for the lower-energy electrons are more prominent, and therefore stronger tails in the phase space distribution are developing during the beam-beam interaction. However, still due to the re-matched beam optics the large majority of the particles can be guided back through the decelerating part of the ERL and an energy recovery efficiency of close to 98% still is obtained.
@@
Figure 5: Visualisation of the beam-beam effect in phase space: in green the ellipse of the ideal configuration is shown. Due to the beam-beam effect, tails in the particle distribution are developing (blue spots) that have to be kept under control for full energy recovery performance. Left: 50 GeV case, right: 20 GeV. The lower energy of the 20 GeV electrons makes them more susceptible to the beam-beam force and the tails are more developed.
@@
$\sigma_{x,e}=\sigma_{x,p}=\sigma_{y,e}=\sigma_{y,p},$
and thus the emittance of the electrons for the 20 GeV case has to be taken into account in a proper manner. In a storage ring the equilibrium emittance is well defined by the emission of synchrotron light. In the case of an ERL however, this equilibrium will never be reached. Ignoring the small effect of quantum emission in return arcs, the electron emittance therefore follows the rule of adiabatic damping $\varepsilon\propto 1/\gamma$. As a consequence, the emittance of the 20 GeV electron beam at the interaction point is considerably larger than the corresponding value at 50 GeV. This fact has to be taken into account to fulfil the matching condition for the transverse sizes between protons and electrons at the IP. In principle a strong reduction of the $\beta$-function would be needed to compensate for the increased beam emittance. However, this is far beyond the capabilities of the mini-beta quadrupoles.
In addition, the conditions of matched beam sizes for the 20 GeV phase-one scenario are obtained with electron beam optics that has been slightly adjusted for a $\beta^{*}$ of 20 cm. Given the emittance and amplitude function of the colliding proton beam that define the optics conditions at the IP, we get a beam size at the $ep$ interaction point of:
@@
For the amplitude function of the electrons, $\beta^{*}=20$ cm we obtain a required absolute (i.e. geometric) electron emittance of $\varepsilon_{0}=5.7\cdot 10^{-4}~{}{\rm mm\cdot mrad,}$. This corresponds to a normalised (energy independent) value of $\varepsilon_{n}=22~{}{\rm mm\cdot mrad}$ and lies well within the specifications of the proposed particle source for the ERL [1]. According to the design parameters of the electron gun a normalised emittance even as low as $\varepsilon_{n}=6~{}{\rm mm}\cdot{\rm mrad}$ is expected (see Tab. 1) which represents a comfortable safety margin for the 20 GeV design.
@@
RMS bunch length mm 3
@@
Uncorrelated energy spread keV 10
Beam polarisation Un-polarised/polarised
In these considerations, the small effect of the emittance increase in the return arcs of the ERL was taken into account. Given the lower beam energy of 20 GeV, the emitted quanta have a much smaller effect, and the resulting emittance dilution remains at a very low level of a few percent.
Beam separation scheme and synchrotron radiation. The beam separation between electrons and protons is provided by the combined effect of the detector weak dipole field and the off-centred quadrupoles of the electron mini-beta concept. Marked in blue in Fig. 2 the combined function quadrupoles of the doublet and the dipole field of the detector represent a bending force with equal and constant bending radius. They provide a smooth separation effect because of the different beam rigidities of electron and proton beams with a minimised critical energy and limited power of the emitted synchrotron light. Applying this scheme to the case of 20 GeV electron energy, results in small values of the emitted light and thus leads to a comfortable situation with $E_{\rm crit}\approx 10~{}{\rm keV}$ and an overall radiated power in the interaction region of well below 1 kW.
Yet, another slightly modified beam separation scheme is possible: the strong dependence of synchrotron radiation on $\gamma$ allows for the 20 GeV case a much stronger bending as the power of the emitted radiation scales as $P_{\gamma}\propto{{\gamma^{4}}/{\rho^{2}}}$ where $\rho$ is the bending radius of the electrons. As a consequence, the applied bending fields can be increased considerably, allowing the electron beam to be guided outside the cryostat of the first proton quadrupole, labelled as Q1 in Fig. 2.
We would like to point out, however, that in any case a certain contribution of the dipole field to the separation is needed to avoid parasitic encounters of the counter-rotating bunches. Following the bunch spacing of the LHC, $\Delta t=25~{}{\rm ns}$, parasitic collisions will occur every 3.75 m and in order to avoid those the beams either have to be brought into collision under a crossing angle (which involves crab cavities to compensate the geometric luminosity loss) or – as foreseen in this scenario – a separation dipole field has to be applied right before and after the collision point, with the total bending power of at least 0.3 Tm.
Cost savings. Following the idea of a design that can be easily extended in energy by staging additional return arcs, the basic layout of the ERL will remain the same. However, starting with one-turn ERL the return arcs 3 to 6 as well as the so-called beam spreaders and re-combiners at the end of the linacs and the bypass that guides the low-energy beam outside the particle detector will not be needed. A cost estimate for the original ERL design results in the budget of magnet and vacuum system of $11~{}{\rm kCHF}$ per metre [6]. As a result, an overall budget reduction of at least $70~{}{\rm MCHF}$ is obtained for the 20 GeV version discussed here. In addition, the lower electron energy allows a much simplified design of the interaction region – in the baseline case, with $\rho\approx 1500~{}{\rm m}$, the synchrotron radiation is relatively weak and does not require installation of sophisticated protection masks and absorbers. In the alternative design with a faster beam separation, using $\rho\approx 120~{}{\rm m}$, the corresponding orbit of the electrons can be guided outside the cryostat of the first superconducting proton quadrupole, located at a 23 m distance from the interaction point, and as a result, the layout of the IR in this case is fully compatible with the existing LHC mini-beta scheme, and a new magnet design for the protons (eventually based on ${\rm Nb}_{3}{\rm Sn}$ technology) is not required for this phase-one scenario. The estimated additional cost savings are about $100~{}{\rm MCHF}$. Last but not least, the operating costs of the 20 GeV ERL will be much smaller with respect to the 50 GeV case because of much lower energy losses due to synchrotron radiation and the absence of the arc 3–6 beamlines.
As discussed in the 2021 CDR revision of the LHeC study [1] the installation of the ERL machine can be made partially in parallel during the Run4 operation period and the detector installation would then take place during the fourth Long Shutdown (LS4) planned in 2035-2036. In total, 7 years are needed to build the LHeC, including permits, engineering studies, tendering and land negotiations, and the final year of connection to HL-LHC [7]. Therefore, given a much simpler design of the phase-one LHeC proposed here and a timely decision on the start of such a project, we believe that the electron-hadron collisions can be available during Run5 of the HL-LHC.
Luminosity reach. The operational scenario for phase-one of the LHeC foresees a truly concurrent running of the $ep$ operation in parallel to the standard HL-LHC proton collisions. As a consequence, the performance of the LHeC depends directly on the HL-LHC design parameters, mainly the frequency of bunch collisions $f_{\rm coll}$, bunch populations $N_{e}$ and $N_{p}$, and the achievable beta function at the IP. In the case of round beams and matched beam sizes, luminosity $\cal L$ is given by:
${\cal L}=\frac{N_{e}N_{p}f_{\rm coll}}{4\pi\varepsilon_{p}\beta^{*}_{p}}~{}\Pi_{i}(\it{H_{i}}),$
@@
where the correction factors for hourglass and pinch effects and for the bunch filling fraction compensate each other, so in a good approximation one expects $\Pi_{i}(H_{i})\approx 1$. It shows that luminosity is basically defined by the achievable amplitude function of the proton beam at the interaction point $\beta^{\ast}_{p}$. Thus, for a given beam size of the protons at the IP, we assume that the optics of the electrons in the ERL beam delivery system will follow, and the limiting factors are the available free aperture of the proton mini-beta quadrupoles and the dynamic effects created by the sextupole fields needed for chromaticity correction.
The values shown in Tab. 2 summarise the situation, with the nominal proton bunch population, $N_{p}=2.2\cdot 10^{11}$ and the electron beam current in the ERL, $I_{e}=60~{}{\rm mA}$, which corresponds to $N_{e}\approx 10^{10}$ electrons per bunch.
@@
Table 2: Luminosity reach in $ep$ collisions for an electron beam energy of 20 GeV and the optimal proton optics definition, assuming a 60 mA electron beam current.
@@
Emittance $\varepsilon$ [$10^{-10}~{}{\rm m}\cdot{\rm rad}$] 3.3 5.7
IP beam size $\sigma^{\ast}$ [$\mu{\rm m}$] 10.7
Luminosity [$10^{33}~{}{\rm cm}^{-2}{\rm s}^{-1}$] $6$

### v2

@@ Abstract
We discuss a concept of a lower-energy version of the Large Hadron-electron Collider (LHeC), delivering electron-hadron collisions concurrently to the hadron-hadron collisions at the high-luminosity LHC at CERN. Assuming the use of a 20 GeV electron Energy Recovery Linac (ERL), we report the results on the optimised beam dynamics, accelerator technologies, and detector constraints required for such a ”phase-one” LHeC. Finally, we also discuss the ERL configurations and the possibility of delivering electron-hadron collisions during the planned Run5 of the LHC, which opens excellent research capabilities – the unique scientific potential of the proposed facility is outlined.
@@ 1 Introduction
The Large Hadron-electron Collider (LHeC) [1, 2] will open new frontiers in high-energy physics as it will become a unique laboratory for studies of deep-inelastic scattering (DIS) and Higgs boson, electroweak, top quark and beyond the Standard Model (BSM) phenomena in a vast range of virtualities $Q^{2}$, from very close to zero up to $10^{6}$ GeV2, and for a much extended range of $x$ – the fractional longitudinal momentum of partons. Recently, an intense electron beam from a novel energy recovery linac (ERL) was proposed to collide with a proton or ion beam from the circular, 27-km circumference High-Luminosity Large Hadron Collider (HL-LHC) at CERN, in the redesigned P2 interaction region capable of hosting electron-hadron or hadron-hadron collisions, concurrently with hadron-hadron collisions in other HL-LHC experiments [3].
In the Conceptual Design Report (CDR) of the LHeC, it is assumed that a three-pass ERL delivers a 50 GeV high-current electron beam with an intensity of up to 20 mA [2], and in the most recent update of the LHeC proposal, its commissioning is proposed several years after the completion of the HL-LHC programme [4]. In this paper, we propose the use of a 20 GeV single-pass ERL with yet a higher beam current of 60 mA, and with a circumference of a third of the LHC’s. This choice of design is motivated by the availability of the technology to operate a single-pass ERL (although presently at lower beam currents) and by the overall simplification, in particular, in the design of the adequate machine-detector interface owing to much lower levels of synchrotron radiation, allowing much earlier commissioning of the high energy electron-hadron collisions at CERN. The electron beam will be accelerated to a full energy of 20 GeV by one pass in two 10 GeV linacs (see Fig. 1), and after collision with the clockwise circulating LHC proton beam, it will be guided back to the linac structures in counter-phase to be completely decelerated within the energy recovery process.
Figure 1: Schematics of the single-pass LHeC layout at the Interaction Point 2 (IP2); with the optional return arcs 363-6 (dashed lines).
The technical feasibility of such a single-pass design has been demonstrated in a number of test facilities and linacs that operate in the range of several mA beam current (see, e.g., [5]). Fig. 2 summarises ERLs that are already operational, in the phase of commissioning or planned.
Figure 2: Different ERL projects worldwide that are already operational or in their commissioning / planning phase. The red and green markers indicate the baseline LHeC design and the ERL configuration described in this paper, respectively. Figure adapted from [5].
Note that the cost of building this ”phase-one” LHeC as well as its running costs are significantly lower than for the final LHeC configuration. Such a staged realisation of the LHeC project provides a huge benefit not only in reducing financial (and personnel) stress but first of all in the availability of concurrent electron-hadron collisions already during Run5 of the current HL-LHC programme. Consequently, apart from the possibility of carrying out a wide range of unique and important studies almost 10 years earlier, the results of this electron-hadron experiment could also be used to improve many important measurements in other (hadron-hadron) experiments at the HL-LHC in a timely manner.
@@ 2 The LHeC interaction region
The fundamental principle of an ERL provides an ideal condition for a staged particle accelerator. Still, a number of items have to be followed up, as beam parameters scale with energy, so while keeping the general layout of the ERL and the geometry of linacs and return arcs, a re-optimisation of the beam dynamics is needed for highest performance of the electron-proton collisions.
Schematically, the concept is shown in Fig. 1 – the geometry of the ERL will remain the same: two superconducting linacs are connected by return arcs on both sides – details about the general layout can be found in [6] and [7]. However, for the phase-one case with reduced electron energy several parts needed to reach higher energies are left out: arcs from 3 to 6 will not be installed, nor will the so-called bypass needed in the final 50 GeV design to guide the beam outside the particle detector, which will only be installed in a later staging step. Beam spreaders and re-combiner modules will also not be present in the phase-one electron beam lattice.
On the other hand, a number of key elements will define the geometry of the ERL from phase one and will remain untouched for all energy steps. The linacs, the interaction region (IR) and the beam separation scheme will be designed to be valid for any operation energy. As described in detail in [6] the proton and electron beams are separated after the interaction point through their different beam rigidities. A combination of a weak dipole field embedded inside the detector and the combined function effect of the off-centred mini-beta quadrupoles of the electron lattice acts as an effective separation scheme. The actual value of the dipole strength and the offset of the quadrupoles were optimised for the lowest synchrotron radiation and thus equal effective bending radii in the three magnets [8], and for the phase-one scenario described here, the electron bending radius of this separation scheme is of about 1500 m.
@@
  • $p$-optics, orbit;
  • beam-beam effect on the proton beam / ditto for the electrons;
  • electron emittance;
  • beam separation scheme and synchrotron radiation;
  • design, work load, and cost optimisation;
  • luminosity reach;
  • polarisation.
Figure 3: The interaction region of electron and proton beams: in black the mini-beta focusing scheme of the electrons is shown, including the detector dipole field that is part of the beam separation. Both are embedded within the free space of the HL-LHC proton inner triplet lattice, marked in blue and red. Figure adapted from [9].
The interaction region of the $ep$ collisions is shown schematically in Fig. 3. Given the comparatively low energy of the electrons, a gradient of approximately g=20g=20~T/m and a length of 1.8 m are sufficient to meet the optics requirements of the electron quadrupoles. These magnets can therefore be made very compact, and the dimensions of the electron mini-beta insertions have been chosen so that they will fit easily in the free space of the IR. Only the first quadrupole of the proton lattice, “Q1” in the figure, will have to be replaced by a new design that allows a field free region for the electrons as their separation is not large enough yet to be guided completely outside the cryostat. For details of the technical design of this new magnet we refer to [2]. With the separation scheme explained here, head-on collisions between the proton and electron beam are foreseen, to avoid luminosity reduction because of the geometric loss factor from an eventual crossing angle.
[…]
Figure 4: Local distortion of the proton optics due to the influence of the electron quadrupoles before (red) and after a local compensation of the effect (blue) for the case of a 20 GeV electron energy. Figure extracted from [9].
Beyond the optics distortion from the electron gradient fields that have a certain influence on the proton optics, the effect of the beam-beam interaction has been studied for both beams. As expected, due to the limited bunch current of the electrons, the proton beam-beam effect has only a very marginal impact on the proton emittance and stability. In the end, the electron bunch intensity is in the order of one percent of the proton bunch population, which leads to a very limited beam-beam force compared to the other LHC proton-proton interaction points. The linear part of the beam-beam interaction as well as the non-linear effect has been studied and long-range forces are not present: due to the fast beam separation scheme, the beam separation between electrons and protons is already well beyond 20 $\sigma$ beam size, and thus – based on LHC operation experience – long-range effects can be neglected. Studies on long-term stability, including orbit fluctuations of the electron beam, are ongoing and might support the need for a fast acting correction system.
@@
where we refer here to the effect on the proton beam; accordingly $N_{e}$ is the electron bunch population, $r_{0}$ the classical proton radius, $\gamma$ is the electron Lorentz factor and $\beta^{\ast}$ and $\sigma_{x,y}$ are the amplitude function and transverse beam sizes at the Interaction Point (IP). We would like to point out that in this approximation the tune-shift does not depend on the $\beta$-function. However, it depends on the energy of the beam considered. For the design case of the 50 GeV electron and 7 TeV proton beams, we estimate the effect using a realistic MAD-X modelling of the HL-LHC lattice including the electron quadrupoles in the IR, and the remaining distortions are compensated; see Fig. 4. Due to the low energy of the electrons – compared to the 7 TeV of the proton beam – the insertion quadrupoles of the electron lattice have only a marginal effect on the proton beam: the resulting beam-beam tune-shift of $\Delta Q_{x,y}\approx 3.8\cdot 10^{-4}$ is nearly two orders of magnitude smaller than the effect of the proton-proton interactions for the HL-LHC operation where we obtain $\Delta Q_{x,y}\approx-8.6\cdot 10^{-3}$ per IP and thus the additional distortion due to $ep$ collisions only has a marginal effect. The distortion of the proton beta function obtained is only $1.8\%$ – without correction. It is corrected nevertheless by slightly modifying the gradients of the neighbouring p-quads (a so-called local rematch of the beam optics).
@@
Beam-beam effect on the electron beam. The situation changes considerably for the electrons. Due to the lower beam energy the electrons are more sensitive, and especially the effect of the non-linear terms has to be minimised. In Fig. 5, we compute the beam size of the electron beam as a function of the position $s$ along the beam orbit with $s=0$ referring to the IP. In black, the theoretical beam size is shown as it is obtained without beam-beam forces. The strong additional focusing due to the space-charge forces of the proton bunch is also shown in blue. The green dashed line quantifies the mismatch with the design linear optics [10]. Here we would like to point out that the dominant parameter, the proton bunch population, refers to the ultimate value of $N_{p}=2.2\cdot 10^{11}$ as foreseen for the HL-LHC luminosity upgrade. Under the influence of this additional focusing – called “pinch effect” – the electron beam reaches a smaller effective beam size which is, however, to a large extent compensated by the so-called hourglass effect in the vicinity of the IR. However, the optics distortion has to be corrected in order to obtain well-matched beam parameters before entering the return arc and the decelerating process of the ERL. This is also highlighted in Fig. 5 where, using the same modelling of the modified HL-LHC lattice introduced above, the black and blue curves are shown to be congruent before and after the interaction region.
@@
Figure 5: Effect of the space charge force on the 20 GeV electron beam, or “beam-beam force”. The beam size is reduced (leading to an additional enhancement of the luminosity, pinch effect) and re-matched to the ideal conditions before entering the return arc for energy recovery.
@@
While this fact represents the prerequisite for a successful energy recovery, the non-linear terms of the beam-beam interaction are better visible in the phase-space representation. We compare the situation for the two electron energies considered here, 50 GeV and 20 GeV, in Fig. 6. As expected, the beam-beam effects for the lower-energy electrons are more prominent, and therefore stronger tails in the phase space distribution are developing during the beam-beam interaction. However, still due to the re-matched beam optics the large majority of the particles can be guided back through the decelerating part of the ERL and an energy recovery efficiency of close to 98% still is obtained.
@@
Figure 6: Visualisation of the simulated beam-beam effect in phase space: in green the ellipse of the ideal configuration is shown. Due to the beam-beam effect, tails in the particle distribution are developing (blue spots) that have to be kept under control for full energy recovery performance. Left: 50 GeV case, right: 20 GeV. The lower energy of the 20 GeV electrons makes them more susceptible to the beam-beam force and the tails are more developed.
@@
$\sigma_{x,e}=\sigma_{x,p}=\sigma_{y,e}=\sigma_{y,p}.$
Thus, the emittance of the electrons for the 20 GeV case has to be taken into account in a proper manner. In a storage ring the equilibrium emittance is well defined by the emission of synchrotron light. In the case of an ERL however, this equilibrium is never reached. Ignoring the small effect of quantum emission in return arcs, the electron emittance therefore follows the rule of adiabatic damping $\varepsilon\propto 1/\gamma$. As a consequence, the emittance of the 20 GeV electron beam at the interaction point is considerably larger than the corresponding value at 50 GeV. This fact has to be taken into account to fulfil the matching condition for the transverse sizes between protons and electrons at the IP. In principle a strong reduction of the $\beta$-function would be needed to compensate for the increased beam emittance. However, this is far beyond the capabilities of the mini-beta quadrupoles.
In addition, the conditions of matched beam sizes for the 20 GeV phase-one scenario are obtained with electron beam optics we slightly adjusted from the final LHeC configuration for a $\beta^{*}$ of 20 cm. Given the emittance and amplitude function of the colliding proton beam that define the optics conditions at the IP, we get a beam size at the $ep$ interaction point of:
@@
For the amplitude function of the electrons, $\beta^{\ast}=20$ cm we obtain a required absolute (i.e. geometric) electron emittance of $\varepsilon_{0}=5.7\cdot 10^{-4}~{\rm mm\cdot mrad}$. This corresponds to a normalised (energy-independent) value of $\varepsilon_{n}=22~{\rm mm\cdot mrad}$ and lies well within the specifications of the proposed particle source for the ERL [2]. According to the design parameters of the electron gun a normalised emittance even as low as $\varepsilon_{n}=6~{\rm mm}\cdot{\rm mrad}$ is expected (see Tab. 1) which represents a comfortable safety margin for the 20 GeV design.
@@
RMS bunch length $m$ 3
Normalised transverse emittance $m\cdotmrad$ < 6
Uncorrelated energy spread keV 10
Beam polarisation Un-polarised/longitudinally polarised ($\pm 80\%$)
In these considerations, the small effect of the emittance increase in the return arcs of the ERL was taken into account. Given the lower beam energy of 20 GeV, the emitted quanta have a much smaller effect, and the resulting emittance dilution remains at a very low level of a few percent. Other effects might influence the quality of the electron beam: the effect of CSR (coherent synchrotron radiation) has been studied in the context of front-to-end simulations that have been performed for the ERL. Given the parameter choice of the machine no significant influence has been found. Beam-Breakup (BBU) as well as micro bunching instabilities on the other side are items still on the to-do-list towards a technical design report as well as tolerance considerations for the beam-beam effect between the non-colliding proton beam and the electrons, for example. They will be investigated in the near future.
Beam separation scheme and synchrotron radiation. The beam separation between electrons and protons is provided by the combined effect of the detector weak dipole field and the off-centred quadrupoles of the electron mini-beta concept. Marked in grey in Fig. 3 the combined function quadrupoles of the doublet and the dipole field of the detector represent a bending force with equal and constant bending radius. They provide a smooth separation effect because of the different beam rigidities of electron and proton beams with a minimised critical energy and limited power of the emitted synchrotron light. Applying this scheme to the case of 20 GeV electron energy results in $E_{\rm crit}\approx 15-20~{\rm keV}$ and an overall radiated power in the interaction region of well below 1 kW. Optimisation work is performed modelling the electron lattice including the first HL-LHC proton quadrupoles close to the interaction region, and using version 1.7.7 of the beam delivery simulation toolkit (BDSIM) [11], With a reduced region of interest centred on the first elements around IP2 (including the first proton quadrupoles), it assumes two scenarios for the free drift length: $L^{\ast}=21$, and $15~{\rm m}$. Although the reduction of the radiated power is clear between these two values of the $L^{\ast}$ for the 50 GeV baseline, it is observed to be negligible for the 20 GeV option.
We would like to point out, however, that in any case a certain contribution of the dipole field to the separation is needed to avoid parasitic encounters of the counter-rotating bunches. Following the bunch spacing of the LHC, $\Delta t=25~{\rm ns}$, parasitic collisions will occur every 3.75 m and in order to avoid those the beams either have to be brought into collision under a crossing angle (which involves crab cavities to compensate for the geometric luminosity loss) or – as foreseen in this scenario – a separation dipole field has to be applied right before and after the collision point, with the total bending power of at least 0.3 Tm.
Design, work load, and cost optimisation. Following the idea of a design that can be easily extended in energy by staging additional return arcs, the basic layout of the ERL will remain the same. Starting with one-pass ERL the return arcs from 3 to 6 as well as the so-called beam spreaders and re-combiners at the end of the linacs and the bypass that guides the low-energy beam outside the particle detector will not be needed. An estimate of the material cost for the original ERL design resulted in a magnet and vacuum system budget of $11~{\rm kCHF}$ per metre [12]. As a result, an overall budget reduction of at least $70~{\rm MCHF}$ is obtained for the 20 GeV version discussed here. In addition, the work load and associated personnel costs are significantly reduced. Moreover, the lower electron energy allows a much simplified design of the interaction region – in the baseline case, with $\rho\approx 1500~{\rm m}$, the synchrotron radiation is relatively weak and does not require installation of sophisticated protection masks and absorbers. Finally, as discussed in the following, detector costs are minimised by joining forces with the ALICE 3 project. Last but not least, the operating costs of the 20 GeV ERL will be much smaller with respect to the 50 GeV case because of much lower energy losses due to synchrotron radiation and the absence of the arc 3–6 beamlines.
As discussed in the 2021 CDR revision of the LHeC study [2] the installation of the ERL machine can be made partially in parallel during the Run4 operation period and the detector installation would then take place during the fourth Long Shutdown (LS4) planned in 2035–6 [13]. In total, 7 years are needed to build the LHeC, including permits, engineering studies, tendering and land negotiations, and the final year of connection to HL-LHC [4]. Therefore, given a much simpler design of the phase-one LHeC proposed here and a timely decision on the start of such a project, we believe that the electron-hadron collisions can be available during Run5 of the HL-LHC which will start, according to the present planning, in 2036.
Last but not least, on the basis of the experience acquired during the construction and commissioning of the phase-one LHeC, the design of the final LHeC will be optimised. Not only at a very technical level, but also in terms of the optimal electron beam parameters and configuration – for example, by making a well-founded choice between two alternatives: A three-pass ERL setup, providing the highest energy but at the same time a somewhat compromised luminosity, against a two-pass ERL configuration resulting in a lower energy but the highest luminosity.
@@
${\cal L}\simeq\frac{1}{4\pi}\frac{N_{e}N_{p}}{\varepsilon_{p}\beta^{\ast}_{p}}f_{\rm coll}.$
It should be mentioned that, given the parameters of the LHeC, the usual luminosity reduction factors do not play a significant role – in particular, head-on collisions between electrons and protons are foreseen at the LHeC. Thus, the luminosity is basically defined by the achievable amplitude function of the proton beam at the interaction point $\beta^{\ast}_{p}$ which is the leading parameter: in order to achieve highest luminosity, the optics design of the protons in the $ep$ interaction region pushes for minimum possible values of $\beta^{\ast}_{p}$, limited by aperture requirements and chromatic budget. The $\beta^{\ast}$ of the electrons on the other side can be chosen in a wide range, due to the flexibility of the linac lattice that – unlike the LHC optics – does not have to observe too many boundary conditions like mini-beta insertions on the other LHC IPs. Accordingly, the combination of $e$-emittance and $\beta^{\ast}$ of the electron optics can easily follow the matched beam conditions of the beam-beam interaction.
@@
As a consequence, for a given beam size of the protons at the IP, we assume that the optics of the electrons in the ERL beam delivery system will follow, and the limiting factors are the available free aperture of the proton mini-beta quadrupoles and the dynamic effects created by the sextupole fields needed for chromaticity correction.
[…]
@@
Table 2: Luminosity reach in epep collisions computed for an electron beam energy of 20 GeV and the optimal proton optics definition, assuming a 60 mA electron beam current.
@@
Emittance $\varepsilon$ $10^{-10}~{\rm m}\cdot{\rm rad}$] 3.3 5.7
IP beam size $\sigma^{\ast}$ $\mu{\rm m}$] 10.7
Bunch population $N_{p,e}$ $10^{10}$] 2.2 0.1
Luminosity $10^{33}~{\rm cm}^{-2}{\rm s}^{-1}$] 66